Suppose we have a small ball, conected by a weightless rigid bar to a rotating point, around which it can easily rotate because of gravity. We hold it vertically and give it a little shake, so that it begins the motion. What is the angle which the bar will form with the initial direction when the vertical component of its speed is maximal?
1
Expert's answer
2020-03-13T11:10:26-0400

In this case "v_{ver}=\sin\alpha\cdot\sqrt{2Rg(1-\cos\alpha)}"


The vertical component is maximal when "\alpha=90°".